Determine the maximum velocity of the resultant of the two waves described by the equation y1=(5.0m)sin((10cm)x-(3.4rad/s)t) and y2=(25m)cos((10cm)x-(3.4rad/s)t) respectively.
Answer #1

Use the properties of interference of waves to find the required maximum speed after finding the amplitude of the resultant wave.yz = A, sin(18e - wt) Yz = A2 cos(k9c-wt) = Az sin(kx-wt+) A, 5m, where A=25m, k = 10cm w = 3-4 radls resu wave So, amplitude

***********************************************************************************************
This concludes the answers. If there is any mistake or omission, let me know immediately and I will fix it....